A 60 kg student slides down a 5.0 metre high slide. At the bottom of the slide , the student is moving at 4.5 m/s (due to friction) . If the slide is a 10.0 Kg layer of steel , by how much will the steel temperature increase?


1
Expert's answer
2021-01-29T01:29:01-0500

Energy lost by the ball will be used in increasing the temperature.


i.e. "\Delta E = \Delta Q"


"m_1gh-\frac{1}{2}m_1v^2 = m_ss\Delta T"


where v is the final velocity of student

h is the height of student

s is the specific heat of steel

"m_1" is the mass of student

"m_s" is the mass of steel

"\Delta T" is the temperature change


Putting values,

"(60 \times 9.8 \times 5) - (\frac{1}{2} \times 60 \times 4.5^2) = 10 \times 420 \times \Delta T"

Solving it we get,

"\Delta T = 0.55 ^\circ C"